OpenStudy (lgbasallote):

Rationalize: \[\huge \frac{\sqrt{4+h} - 2}h\]

OpenStudy (anonymous):

\[\frac{ 4+h-4 }{ h(\sqrt{4+h}+2) }=\frac{ 1 }{\sqrt{4+h}+2 }\] nomarlly limit question

OpenStudy (anonymous):

we multiplied by\[\sqrt{4+h}+2\] both num and denominator

OpenStudy (lgbasallote):

hmmm...isn't the point of rationalization to remove radicals from the denominator?

OpenStudy (swissgirl):

Well then there were no radicals in the denominator

OpenStudy (swissgirl):

So u wld keep it the way it is

OpenStudy (lgbasallote):

really?

OpenStudy (anonymous):

imagine if this question was\[\lim_{h \rightarrow 0}\frac{ \sqrt{4+h}-2 }{ h }\]

OpenStudy (lgbasallote):

??

OpenStudy (anonymous):

you cant just plug h=0 here but in the rationalised one

OpenStudy (swissgirl):

Ohh that is cool :P

OpenStudy (lgbasallote):

limits in an algebra question? that's morbid.....

Parth (parthkohli):

Rationalizing means to remove the radical from the place it is in...

OpenStudy (swissgirl):

hmmm well u can be asked to rationalize the numerator too

OpenStudy (lgbasallote):

now im confused with the contradictions...

OpenStudy (swissgirl):

What are the contradictions?

OpenStudy (lgbasallote):

@swissgirl said don't change...now she says change

OpenStudy (swissgirl):

Just use ur own brain -_-

OpenStudy (swissgirl):

lol

OpenStudy (lgbasallote):

someone's wrong here....wonder who

Parth (parthkohli):

If they ask you to rationalize the fraction, you do these: 1) If the radical is in the denominator — remove it from the denominator. Do not care about the numerator. 2) If the radical is in the numerator — remove it from the numerator. Do not care about the denominator.

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} }=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

rationalised

OpenStudy (lgbasallote):

i know what happens if it's in the denominator

OpenStudy (swissgirl):

trust me lgba knows how to rationalize a numerator or denomanator

OpenStudy (lgbasallote):

the question is if it's in the numerator

Parth (parthkohli):

See how you can't remove the radical from both numerator and denominator?

Parth (parthkohli):

Yes, so move it to the denominator. Do not care about the denominator.

OpenStudy (lgbasallote):

my question is not how to rationalize @swissgirl but if it's suppose to be rationalized

Parth (parthkohli):

You can rationalize it, but mathematicians always love if the radical is in the numerator.

OpenStudy (lgbasallote):

because what i know is that if it's in the numerator, then it's okay

Parth (parthkohli):

It's better to put radicals in the numerator rather than the denominator.

OpenStudy (lgbasallote):

so why put in denominator then?

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }=\frac{ 1 }{ \sqrt{2} }\] these is also called rationalising sometimes it is convinient to write in the dinominatpr eg in the case of a limit

OpenStudy (swissgirl):

Well what i have always learnt was that rationalization=denominator but I have gone online and seen that some do rationalize the numerator too so it all depends what course you are taking and what you see in your the textbook

Parth (parthkohli):

Yup, I saw that you could rationalize the numerator too though it's rare.

OpenStudy (lgbasallote):

when i went online, all i saw were unreliable sources on rationalizing numerators

Parth (parthkohli):

Never rationalize the numerator unless given a question to perform.

OpenStudy (lgbasallote):

\[\lim_{x \rightarrow 0} \sqrt x\]seems better than \[\lim_{x \rightarrow 0} \frac 1{\sqrt x}\] @Jonask

OpenStudy (swissgirl):

These are both reliable sources

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

how ever if you where asked to do the same function using the first prinsiple you will need to rationalise

OpenStudy (lgbasallote):

the links win. can't argue with that.

OpenStudy (anonymous):

\[\frac{ \sqrt{x+h}-\sqrt{x} }{ h }\]ration. \[\frac{ 1 }{ \sqrt{x+h}+\sqrt{x} }\]

OpenStudy (lgbasallote):

square roots in denominators really look weird...